Sue is four years older than Mark. Twice Sue’s age plus 5 times Mark’s age is equal to 64. What are the ages of Sue and Mark?
Answer by richwmiller(17219)

s=m+4
2s+5m=64
2*(m+4)+5m=64
2m+8+5m=64
7m=56
m=8
s=12